OpenStudy (anonymous):

PLEASE HELP ME!!! (2sqrt(3)-1) times (3sqrt(3)-2)

OpenStudy (08surya):

pls draw it and post it

OpenStudy (anonymous):

\[(2\sqrt{3}-1)(3\sqrt{3}-2)\]

OpenStudy (anonymous):

Just simply expand your expression.

OpenStudy (anonymous):

\[(2\sqrt3-1) \times (3\sqrt3-2)\] is this your problem? @rdavis1997

OpenStudy (08surya):

ya multiply each term one by one

OpenStudy (anonymous):

Yes. Sorry guys. Lol, i was doing something.

OpenStudy (anonymous):

@dpasingh @08surya @Noliec

OpenStudy (anonymous):

@rdavis1997 yes.

OpenStudy (anonymous):

Use that (a+b)*(c+d)=a*c+a*d+b*c+b*d

OpenStudy (anonymous):

Also note that, for example, \[\sqrt{2}*\sqrt{2}=2\]

OpenStudy (anonymous):

I need help on understanding. I got an answer but i think its wrong. 20- 7\[\sqrt{3}\]

OpenStudy (anonymous):

The 7 sqrt of 3. sorry

OpenStudy (anonymous):

2sqrt(3)*3sqrt(3) is the same thing as 6*sqrt(3)*sqrt(3)

OpenStudy (anonymous):

6 sqrt(3)

OpenStudy (anonymous):

6*(sqrt(3))^2=6*3

OpenStudy (anonymous):

18

OpenStudy (anonymous):

That was just a stage of the solving; a crucial one, however. :D

OpenStudy (anonymous):

As I said earlier, just expand the terms like you would with any other equation.

OpenStudy (anonymous):

How does that explain if i am right or wrong?

OpenStudy (anonymous):

I have absolutely no idea on what to do.

OpenStudy (anonymous):

Expand!

OpenStudy (anonymous):

I think you confused me a bit more than i already was.

OpenStudy (anonymous):

If I simply give you the solution; you'll be deprived forever off solving it yourself. Meaning you learn immensely much more than you would if I just gave you the solution.

OpenStudy (anonymous):

If you at some stage don't understand what to do, I'll gladly lend a hand.

OpenStudy (anonymous):

i just want you to start me off on a step by step and then ill take the last parts as you back me up.

OpenStudy (anonymous):

Ok, do you know how to expand two terms?

OpenStudy (anonymous):

foil?

OpenStudy (anonymous):

Had to check up what that was, hehe. Yes, foil.

OpenStudy (anonymous):

oh lol nice

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):